(e^x)^x*e^15=e^(8x)
Answer by Alan3354(69443)

(e^x)^x*e^15=e^(8x)
e^(x^2 + 15) = e^(8x)
x^2 + 15 = 8x
x^2 - 8x + 15 = 0
(x-3)(x-5) = 0
x = 3
x = 5